OpenStudy (anonymous):

Help Determine which system below will produce infinitely many solutions. −6x + 3y = 18 4x − 3y = 6 2x + 4y = 24 6x + 12y = 36 3x − y = 14 −9x + 3y = −42 5x + 2y = 13 −x + 4y = −6

OpenStudy (anonymous):

got it C. 3x - y = 14 -9x + 3y = -42

OpenStudy (anonymous):

A:−6x + 3y = 18 4x − 3y = 6 B: 2x + 4y = 24 6x + 12y = 36 C:3x − y = 14 −9x + 3y = −42 D: 5x + 2y = 13 −x + 4y = −6

OpenStudy (anonymous):

Are we sure its C?

OpenStudy (anonymous):

C because 3x-y= 14 multiplied by -3 is -9x+3y= -42 so the equations are equal

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions. @JOHN__22

OpenStudy (anonymous):

@JOHN__22 @JOHN__22